6 Is how many different combinations there are
What is a number that describes how likely it is that a particular event will occur?
The probability.
What is the degree of freedom is the probability is 0.05?
There is no direct relationship between degrees of freedom and probability values.
What is the Greek or Latin word for 'probably'?
The Latin word is probabiliter. The word fortassecan also have this meaning, but is more usually translated "possibly; perhaps".
Full Question:
Among a group of people, 10% are from NYC and 90% are not. All people from NYC are under 18 years old, and 50% of those not from NYC are also under 18 years old. One person is chosen at random, and he or she is under 18. What is the probability that the person chosen is from NYC? Round your answer to two decimal places.
When rolling two dice what is the probability of the sum being a factor of twelve?
That's a very nice problem !
Total number of possible outcomes with two dice = 36.
Factors of twelve = 1, 2, 3, 4, 6, and 12.
Number of ways to roll a '1' = zero
Number of ways to roll a 2 = 1 . .. . (snake-eyes)
Number of ways to roll 3 = 2 . . . . . (1,2 and 2,1)
Number of ways to roll 4 = 3 . . . . . (1,3 / 2,2 and 3/1)
Number of ways to roll 6 = 5 . . . . . (1,5 / 2,4 / 3,3 / 4,2 and 5,1)
Number of ways to roll 12 = 1 . . . . (box-cars)
Total number of successful rolls = 12
Probability = 12/36 = 1/3 = 33 and 1/3 percent.
The probability of drawing a particular color is the number of blocks that color, over the total number of blocks. Since these are disjoint events, the solution is the product of the two probabilities of the draws. So:
P(red, blue) = P(red) * P(blue) = (3/10) * (5/10) = 15/100 or 3/20
To answer this, the total number of questions on the test would need to be known.
What is the probability of getting electric shock?
For mains electricity it is close to 0 if you behave responsibly. For static electricity it depends on the environment that you live in: indoors in cold climates, with synthetic fibres in carpets and dry air, the probability is quite high. However, many people will not register the shock even if there is an electrical discharge.
What is the probability of obtaining at least one 4 in three rolls of a dice?
The probabilty would be 3:18 or 1:6 because you have 3 dice all have 6 sides you do 3*6=18 for the denominator. Now for the numoratoreach die has numbers 1-6 and theres is only for on each dice and there is 3 dice so you would end up having 3 as the numorator. To get one fith you have to simplifie but i will go onto that next time.
1/6*1/2 = 1/12
What is the probability of drawing a diamond in a pack pf 52 cards?
n(S)=52
let A be the event that card drawn is a diamond
then n(A)=13 so,
P(A)=13/52=1/4
If a die is rolled 3 times what is probability or rolling at least one 6?
Probability(at least 1 six) = 1 - probability(no sixes)
= 1 - 5/6 x 5/6 x 5/6
= 1 - 125/216
= 91/216
≈ 0.421 = 42.1%
Nova net answer: They are the same